# JEL Classification: C51, C81, D83, O13, Q12, Q54

### v1

@@ 2 Study Context
Recent data collection efforts, including an RCT (Dar etΒ al., 2013) and the Rice Monitoring Survey (RMS) (Kretzschmar etΒ al., 2018), have provided useful information on STRVs, but they each provide a circumscribed picture of STRVs that is limited by the purpose and methodology of each program. While the RCT has strong internal validity for measuring the impact of STRVs, it has weak external validity as it focused on a small set of villages in a single Indian state and as treated farmers were given the seeds. And, while the RMS provides strong external validity for the breadth of adoption, it has weak internal validity as it started collecting observational data four years after the initial introduction of STRVs. Neither source of data can - nor were they designed to - provide evidence on the long-term, large-scale impacts of IRRIβs (and otherβs) investment in developing STRVs nor BMGFβs investment in producing and disseminating STRVs.
@@
As such, we are operating in a data poor environment. No historical data exist that provide the necessary information: 1) a panel, 2) observations before 2010 and well after 2010 (to allow time for dissemination and adoption), 3) representative at the national level or at least of rice cultivation, and 4) have data on rice varieties grown and yields.555There are a number of nation-wide panel surveys of households in Bangladesh that might appear to be good candidates for filling this data gap. In Online AppendixΒ B we discuss a variety of available data sets and why they are insufficient to answer our research question. And current or future data collection efforts will not be useful as STRVs were introduced over a decade ago, meaning no relevant baseline can be constructed.
@@
As an alternative, we use EO data to reconstruct the past. EO data provide a repeated time series for a given location (panel data), provide observations long before and after 2010, and provide coverage of an entire country. What EO data lack are information of varieties grown, which we can proxy for with administrative data on dissemination efforts, and yields, which we can proxy for with EVI. We can also enrich and validate the EO-based analysis using existing - if incomplete - survey data, such as the RMS, which has data of varieties and yields, the very items that the EO data is missing. Although, there is no way to generate the data one would want if one were planning a prospective impact evaluation of a new agricultural technology, our method provides a way for researchers to answer important questions about the impact of technology adoption in data poor settings.
@@
As a test case for our method, we focus on the adoption of Swarna-Sub1 and subsequent submergence tolerant varieties (which we collectively term STRVs) in flood-prone rice-growing regions of Bangladesh during the Aman season. In the Aman season, farmers face more crop loss due to natural disasters like flood (Bairagi etΒ al., 2021; Khan and Roy, 2020). The intuition behind our identification strategy is simple: for a given rice-growing location, if no flooding occurs, EVI provides a consistent signal of crop growth up until harvest. For that location, prior to 2010, flooding of amount would cause a consistent signal response in which EVI values abruptly drop. Post introduction of STRVs, if there is no flooding, there would be no change in the EVI signal relative to before 2010. And, if there is flooding of amount and no adoption of STRVs, there would again be no change in the signal response relative to before 2010. But, if there is a change in the signal response to flooding, the only explanation would be the adoption of STRVs, as no other technology to reduce the negative impact of flooding on rice was concurrently introduced.
@@ 3 Data
Data availability is the key challenge to establishing the long-term, large-scale impacts of agricultural technology in international development. As Burke etΒ al. (2021) point out, the relative infrequency of in-person surveys in many developing countries can potentially be overcome using remotely sensed EO. In the case of STRVs in South Asia, where no comprehensive baseline data was collected, EO data allows us to look back into the past and reconstruct outcomes before and after the technology was disseminated.
In this study we combine data from various sources to provide evidence on the impact of STRV adoption from a variety of angles. For our national-level analysis, we combine data from MODIS with two types of ground truth data to build rice area maps. We also use MODIS to extract the enhanced vegetation index (EVI), which is our primary proxy for rice yields. To this we layer on flood maps built through a fusion model developed by Giezendanner etΒ al. (2023) that combines data from Sentinel-1 and MODIS. Finally, we combine the EO data with administrative data at the district-level on the production and dissemination of STRV seeds by both public and private entities. For our household-level analysis, we combine three rounds of panel data collected as part of the Rice Monitoring Survey (RMS) with historic flood data from the fusion model. We supplement this with cross-sectional, plot-level data also drawn from the RMS.
The combination of these various sources of data provide us with a picture of the impacts of STRV adoption at a variety of spatial and temporal resolutions. While we view no single source of data or econometric approach to analysis as definitive in establishing the causal impacts of adoption, the sum of these various sources and methods provides a clear picture and a robust understanding of the impacts of this technology.
3.1 Remote Sensing EO Data
@@
Our construction of the EO data set starts by identifying where rice was cultivated in Bangladesh. From the time series of rice maps, we extract EVI values. We then construct flood maps for areas where rice was grown. All of this pixel-level data is aggregated to the district-level (there are 64 districts in Bangladesh) and then combined with administrative data on seed dissemination. In this section, we provide a non-technical, intuitive summary of the methods used to generate the EO data. Online AppendixΒ C provides the technical details sufficient to implement our procedures in alternative contexts.
3.1.1 Rice Area Mapping
@@
As early as Xiao etΒ al. (2006), rice pixels have been identified by detecting surface water followed by a sudden growth of vegetation. The detection of surface water is critical for identification as it represents agronomic flooding, which is unique in rice cultivation and the most common practice for cultivating rice in Bangladesh.666In recent years there has been a push by IRRI and other development organizations to promote alternate wetting and drying (AWD) and/or direct seeded rice (DSR), both of which would substantially reduce or completely eliminate the need for agronomic flooding. However, promotion of both methods has mainly focused on use during the Boro (dry) season when groundwater irrigation is necessary for cultivation. To date, there is a dearth of reliable data to measure adoption of either technology in Bangladesh. See Alauddin etΒ al. (2020) for a recent overview of AWD in Bangladesh. In early applications of remote sensing, which relied on handcrafted, indice-based features, agronomic flooding is signaled when the value of vegetation indices (e.g., the Enhanced Vegetation Index [EVI]), are lower than the value of water indices (e.g., Land Surface Water index [LSWI]). The sudden growth of vegetation is determined if vegetation index reaches half of the maximum value within 40 days immediately after agronomic flooding.
@@
If our goal was to simply construct rice area maps that were contemporaneous with the RMS household data collection efforts (2014, 2017, and 2022), then we could follow a process similar to Lobell etΒ al. (2020) in validating the maps. This would involve taking GPS measures of the area of a number of rice and non-rice plots and using them as ground truth data for training a machine learning classification algorithm. The algorithm could then use the ground truth training data to predict rice and non-rice areas for the rest of the country for that year. However, our goal is to build rice area maps for the 20 year period of 2002-2021, providing us with a large in both the pre- and post-STRV release period. No traditional ground truth data exists for predicting where rice is grown during years when no RMS data was collected.
@@
To address this data deficit, we take two approaches designed to complement one other. Our first approach involves collecting traditional ground truth data. In 2021, we conducted a small survey of RMS households to collect plot GPS coordinates, crop varieties, and planting dates . There was no socioeconomic component of this survey as it was solely focused on building a ground truth data set for training. This effort was augmented in 2022 by the third round of RMS data collection, which included household socioeconomic information plus plot production information and GPS locations . All plot location data was collected following guidance in Azzari etΒ al. (2021) and allows for creating rice area maps in 2021 and 2022 following standard procedures as in Lobell etΒ al. (2020).
@@
Our second approach re-constructs βground truthβ through visual inspection of high resolution Google Earth images. This visual reconstruction using optical imagery is commonly used in the remote sensing literature and is particularly useful for capturing difficult to measure activities, like illegal fishing (Park etΒ al., 2020). In our case, it is difficult to rice cultivation, spanning back two decades. To build our historical ground truth data set, we selected three districts in Bangladesh (Barisal, Kurigam, and Rajshahi) that experience different hydrological characteristics and thus different distributions of riceland. We derived grids from the MODIS pixels and selected a stratified random sample of 150 points (75 points each for rice and non-rice areas) in each district for a total of 450 points (see FigureΒ 2). This was completed for nine years for which Google Earth imagery exists: 2002, 04, 06, 09, 15, 16, 18, 19 and 21. The MODIS grids were overlaid on Google Earth imagery as the basis for visually determining the land cover of each grid for the particular season being validated. MODIS pixels were categorized as rice if, in the Google Earth imagery, more than of the grid was visually determined to be rice. Pixels were categorized as non-rice otherwise.
@@
We then classify rice areas for all years for the entire country by training a random forest (RF) algorithm on the ground truth data using a leave-one-out cross validation scheme. The input data to the model is a combination of surface reflectance and vegetation indices (specifically EVI) extracted from the MODIS eight- and sixteen-day composites at 250m resolution, as well as topographical indicators (elevation and slope, derived from the Digital Elevation Model (DEM) product FABDEMΒ (Hawker etΒ al., 2022)). The RF model consists of 1,000 trees each with five leaves. We use a standard leave-one-out cross validation scheme in which we rotate the districts used as training, leaving one district out each time. This ensures generalizability of the model over space. To ensure generalizability over time, we reserve 2020 as an out-of-sample test set. Our model is accurate at predicting rice and non-rice pixels in 2020 of the time.777Online AppendixΒ C.1 contains more details on the methods, including results of accuracy tests for the model. The end results is a time series of rice and non-rice pixels for the entire country for all years spanning 2002 to 2022.
3.1.2 Flood Area Mapping
@@
As with the rice maps, building reliable flood maps that look into the past raises data availability challenges. Until recently, two common approaches to using EO data for generating flood maps was either to rely on the historic flood data base curated by the Dartmouth Flood Observatory (Brakenridge, nd) or to use MODIS to construct the Modified Normalized Difference Water Index (MNDWI) (Guiteras etΒ al., 2015; Chen etΒ al., 2017; Chen and Mueller, 2018).
@@
The archive from the Dartmouth Flood Observatory, which extends back to 1985, has been highly cited but suffers from several well-know data deficiencies. Most importantly, it only includes large flood events that members of the observatory were able to identify in news stories and governmental sources. Many small floods, that might destroy crops for some farmers but result in no deaths and generally do not make news, are not covered in the archive. Recently, Tellman etΒ al. (2021) develop the Global Flood Database (GFD), which uses 12,719 scenes from Terra and Aqua MODIS sensors to produce daily images of flooding at 250m resolution from 2000 to 2018. The GFD builds on the flood event reported in the Dartmouth data and is validated on 30m resolution Landsat data. While the GFD is based on MODIS 250m, it provides only a binary flood/not-flood indicator at that resolution.
@@
In terms of using MNDWI and MODIS, the MNDWI is a threshold method for identifying flooding developed by Xu (2006). Water and non-water are identified based on surface reflectance and a pixel is defined as water if MNDWI is above zero. This handcrafted feature has been popular because it requires no ground truth data and is simple to construct. However, as Burke etΒ al. (2021) point out, methods that combine EO data with ground truth data and methods from deep learning can make substantial improvements to accuracy relative to previous approaches, like EO-only determined thresholds. Furthermore, using MODIS or any optical sensor only to map floods can cause large underestimations in inundation due to cloud cover.
Sentinel-1 provides a numbers of improvements over MODIS and the Dartmouth data that address these two issues (low resolution and underestimation). Specifically, Sentinel-1 comes at a spatial resolution of 10m and uses a synthetic aperture radar (SAR) sensor, which allows it to penetrateins cloud cover. The limitation to Sentinel-1, for our purposes, is that data exists only back to 2017. To overcome the lack of historic Sentinel-1 data, we train a Convolutional Neural β Long Short-Term Memory Network (CNNβLSTM) on Sentinel-1 fractional flood data to predict historic MODIS satellite data and then project back into the past, prior to the launch of Sentinel-1. The method is described in detail in Giezendanner etΒ al. (2023) and the workflow is summarized in FigureΒ 2.
@@
To detect water in Sentinel-1 data we use a dynamic thresholding algorithm tailored to Bangladesh by Thomas etΒ al. (2023), which produces a binary flood indicator for the 10m pixel. This is upscaled to create a nearly continuous fractional index at 500m resolution, which matches the resolution in MODISβs eight-day composite images. We then feed this Sentinel-1 derived fractional index, along with hydrologically relevant indices (elevation, slope, and height above nearest drainage), and MODIS 500m eight-day composite data into a CNN-LSTM to predict fractional flood area back to 2002. Relative to other deep learning frameworks, the CNN-LSTM is able to exploit the spatial and temporal structure of the data. The data (features) from a single time step are feed into a first CNN which produces a single pixel of spatially contextualized data. This is repeated for each time step to produce a time series of length for that pixel. We then feed a time series of length into the LSTM to produce a single temporal value. The output of the LSTM is then combined with the output of the CNN for time and feed into a second CNN to produce a fractional value for the extent of flooding in a given pixel for a given time step. As with the RF model for rice area, the CNN-LSTM is trained using leave-one-out cross validation, where a time step is withheld for validation one at a time. Saunders etΒ al. (2023) shows that this method produces accuracy levels comparable to existing flood detection methods using MODIS data and previous methods like MNDWI. And, in our case, our model is accurate of the time.888Online AppendixΒ C.2 provides more details about the fusion model and tests of the model against traditional methods. The result is weekly data on the fraction of a 500m pixel that was flood, covering all of Bangladesh from 2002 to 2022.
3.1.3 Creation of District-Level EO Data Set

### v2

@@ 2.3 A Data Poor Setting
Data exists on both the adoption and impact of STRVs, however no data exists that would allow us to study the long-term, large-scale impact of adopting STRVs on rice production. Previous data collection efforts, including an RCT in India (Dar etΒ al., 2013) and a nationally representative survey in Bangladesh (Kretzschmar etΒ al., 2018), provided useful information on STRVs. But, they each provide a circumscribed picture that is limited by the purpose and methodology of each program. While the RCT has strong internal validity for measuring the impact of STRVs, it has weak external validity as it focused on a small set of villages in a single Indian state and as treated farmers were given the seeds. And, while the survey provides strong external validity for the breadth of adoption across Bangladesh, it has weak internal validity as it started collecting observational data four years after the initial introduction of STRVs.
@@
With all of this taken together, it is clear that we are operating in a data poor environment. No historical data exist that provide the necessary information: 1) a panel, 2) observations before 2010 and well after 2010 (to allow time for dissemination and adoption), 3) representativeness at the national level or at least representativeness of rice cultivation, and 4) precise data on length of flooding, yields, and rice varieties grown. And current or future data collection efforts will not be useful as STRVs were introduced over a decade ago, meaning no relevant baseline can be constructed.555There are a number of nation-wide panel surveys of households in Bangladesh that might appear to be good candidates for filling this data gap. In Online AppendixΒ C we discuss a variety of available datasets and why they are insufficient to answer our research question.
@@
As an alternative, we use EO data to reconstruct the past. EO data provides a repeated time series for a given location (panel data), provides observations long before and after 2010, provides coverage of an entire country, and provides spectral information that can be used to identify floods and yields. What EO data lacks is information of varieties grown, which we can proxy for with administrative data on dissemination efforts. We can also enrich and validate the EO-based analysis using existing, if incomplete, survey data. Although, there is no way to generate the data one would want if one were planning a prospective impact evaluation of a new agricultural technology, our method provides a way for researchers to answer important questions about the impact of technology adoption in data poor settings, particularly the past.
@@
As a test case for our method, we focus on the adoption of Swarna-Sub1 and subsequent submergence tolerant varieties (which we collectively term STRVs) in flood-prone rice-growing regions of Bangladesh during the Aman (wet) season. In the Aman season, farmers face more crop loss due to natural disasters like flooding (Bairagi etΒ al., 2021; Khan and Roy, 2020). The intuition behind our identification strategy is simple: for a given rice-growing location, if no flooding occurs, EVI provides a consistent signal of crop growth up until harvest. For that location, prior to 2010, a Goldilocks flood of amount would cause a consistent signal response in which EVI values abruptly drop. After the introduction of STRVs, if there is no flooding, there would be no change in the EVI signal in that location relative to before 2010. And, if there is flooding of amount and no adoption of STRVs, there would again be no change in the signal response in that location relative to before 2010. Further, if there is flooding and adoption of STRVs, but that flooding is not , meaning it is outside the relevant flood window, there would also be no change in the signal response. Only if there is the βjust rightβ amount of flooding and adoption of STRVs would we expect a change in the signal response. If we see such a change in the signal response in a given location to a Goldilocks flood post-2010 relative to a Goldilocks flood pre-2010, the only explanation would be the adoption of STRVs, as no other technology to reduce the negative impact of flooding on rice was concurrently introduced.
@@ 3 Data
Data availability is the key challenge to establishing the long-term, large-scale impacts of agricultural technology in international development. As Burke etΒ al. (2021) point out, the relative infrequency of in-person surveys in many developing countries can potentially be overcome using remotely sensed EO data. In the case of STRVs in South Asia, where no comprehensive baseline data was collected, EO data allows us to look back into the past and reconstruct outcomes before and after the technology was disseminated.
@@
Construction of our EO dataset starts by identifying where rice was cultivated in Bangladesh. From the time series of rice maps, we extract EVI values. We then construct flood maps for areas where rice was grown. All of this pixel-level data is aggregated to the district-level (there are 64 districts in Bangladesh) and then combined with district-level administrative data on seed dissemination. In this section, we provide a non-technical, intuitive summary of the methods used to generate the EO data.666Online AppendixΒ D provides the technical details sufficient to implement our procedures in alternative contexts.
3.1 Rice Area Mapping
@@
As early as Xiao etΒ al. (2006), rice pixels have been identified by detecting surface water followed by a sudden growth of vegetation. The detection of surface water is critical for identification as it represents agronomic flooding, which is unique in rice cultivation and the most common practice for cultivating rice in Bangladesh.777In recent years there has been a push by IRRI and other development organizations to promote alternate wetting and drying (AWD) and/or direct seeded rice (DSR), both of which would substantially reduce or completely eliminate the need for agronomic flooding. However, promotion of both methods has mainly focused on use during the Boro (dry) season when groundwater irrigation is necessary for cultivation. To date, there is a dearth of reliable data to measure adoption of either technology in Bangladesh. See Alauddin etΒ al. (2020) for a recent overview of AWD in Bangladesh. In early applications of remote sensing, which relied on handcrafted, indice-based features, agronomic flooding is signaled when the value of vegetation indices (e.g., the Enhanced Vegetation Index [EVI]), are lower than the value of water indices (e.g., Land Surface Water Index [LSWI]). The sudden growth of vegetation is determined if the vegetation index reaches half of the maximum value within 40 days immediately after agronomic flooding.
@@
If our goal was to simply construct contemporaneous rice area maps, then we could follow more recent methods similar to Lobell etΒ al. (2020). This would involve taking GPS measures of the area of a number of rice and non-rice plots and using them as ground truth data for training a machine learning classification algorithm. The algorithm could then use the ground truth training data to predict rice and non-rice areas for the rest of the country for that year. However, our goal is to build rice area maps for the 20 year period of 2002-2021, providing us with a large in both the pre- and post-STRV release period. No traditional ground truth data exists for predicting where rice is grown 20 year ago.
@@
To address this data deficit, we take two approaches designed to complement one other. Our first approach involves collecting traditional ground truth data. In 2021, we conducted a small survey to collect ground truth data from households, including plot GPS coordinates, crop varieties, and planting dates . This effort was augmented in 2022 by a third round of the Rice Monitoring Survey (RMS), which included household socioeconomic information plus plot production information and GPS locations . All plot location data was collected following guidance in Azzari etΒ al. (2021) and allows for creating rice area maps in 2021 and 2022 following standard procedures as in Lobell etΒ al. (2020).
@@
Our second approach re-constructs βground truthβ through visual inspection of high resolution Google Earth images. This visual reconstruction using optical imagery is particularly useful for capturing difficult to measure activities, like illegal fishing (Park etΒ al., 2020). In our case, it is difficult to capture rice cultivation spanning back two decades. To build our historical reference dataset, we selected three districts in Bangladesh (Barisal, Kurigam, and Rajshahi) that experience different hydrological characteristics and thus different distributions of riceland. We derived grids from the MODIS pixels and selected a stratified random sample of 150 points (75 points each for rice and non-rice areas) in each district for a total of 450 points (see Panel A of FigureΒ 3). This was completed for the nine years for which Google Earth imagery exists: 2002, β04, β06, β09, β15, β16, β18, β19 and β21. The MODIS grids were overlaid on the Google Earth imagery as the basis for visually determining the land cover of each grid for the particular season being validated. MODIS pixels were categorized as rice if, in the Google Earth imagery, more than of the grid was visually determined to be rice. Pixels were categorized as non-rice otherwise.
@@
We then classify MODIS pixels for all years for the entire country as rice/not-rice by training a random forest (RF) algorithm on the ground truth data. The input data to the model is a combination of surface reflectance and EVI extracted from the MODIS eight- and sixteen-day composites at 500m resolution, as well as topographical indicators including elevation and slope, derived from the Digital Elevation Model (DEM) product FABDEMΒ (Hawker etΒ al., 2022). The RF model consists of 1,000 trees each with five leaves. We use a standard leave-one-out cross validation scheme in which we rotate the districts used as training, leaving one district out each time. This ensures generalizability of the model over space. To ensure generalizability over time, we reserve 2020 as an out-of-sample test set. Our model is accurate at predicting rice and non-rice pixels in 2020 of the time.888Online AppendixΒ D.1 contains more details on the methods, including results of accuracy tests for the model. The end result is a time series of rice and not-rice pixels for the entire country for all years spanning 2002 to 2022.
3.2 Flood Area Mapping
@@
As with the rice maps, building reliable flood maps that look into the past raises data availability challenges. Until recently, two common approaches to using EO data for generating flood maps before 2015 (when public radar imagery and higher spatial resolution data from Sentinel-1/2 became available) was either to rely on the historic flood data base curated by the Dartmouth Flood Observatory (DFO) (Brakenridge, nd) or to use MODIS to construct the Modified Normalized Difference Water Index (MNDWI) (Guiteras etΒ al., 2015; Chen etΒ al., 2017; Chen and Mueller, 2018).
@@
The archive from the DFO, which extends back to 1985, has been highly cited but suffers from several well-known data deficiencies. Most importantly, it only includes large flood events that members of the observatory were able to identify in news stories and governmental sources. Many small floods, that might destroy crops for some farmers but result in no deaths and generally do not make news, are not covered in the archive. Recently, Tellman etΒ al. (2021) develop the Global Flood Database (GFD), which uses 12,719 scenes from Terra and Aqua MODIS sensors to produce daily images of flooding at 250m resolution from 2000 to 2018. The GFD builds on the flood events reported in the DFO data and is validated on 30m resolution Landsat data. While the GFD is based on MODIS 250m, it provides only a binary flood and not-flood indicator at that resolution and only for events cataloged by the DFO.
@@
In terms of using MNDWI and MODIS, the MNDWI is a threshold method for identifying flooding developed by Xu (2006). Water and non-water are identified based on surface reflectance and a pixel is defined as water if MNDWI is above zero. As with the handcrafted feature for identifying rice, this handcrafted flood feature has been popular as it requires no ground truth data and is simple to construct. However, as Burke etΒ al. (2021) point out, methods that combine EO data with ground truth data and methods from deep learning can make substantial improvements to accuracy relative to previous approaches, like EO-only determined thresholds. Furthermore, using only MODIS, or any optical sensor, to map floods can cause large underestimations in inundation due to cloud cover. Thus, current remote sensing methods to measure flooding are likely to be highly imprecise in predicting the vary variable critical to overcoming the Goldilocks Problem in STRVs.
Since 2017, Sentinel-1 has provided a numbers of improvements over MODIS and the DFO data and it addresses the issues of low resolution and underestimation. Specifically, Sentinel-1 comes at a spatial resolution of 10m and uses a synthetic aperture radar (SAR) sensor, which allows it to penetrate cloud cover. The limitation to Sentinel-1, for our purposes, is its relative recency. To overcome the lack of historic Sentinel-1 data, we train a Convolutional Neural Network β Long Short-Term Memory (CNNβLSTM) on Sentinel-1 fractional flood data to predict historic MODIS satellite data and then project back into the past, prior to the launch of Sentinel-1. The method is described in detail in Giezendanner etΒ al. (2023).999Online AppendixΒ D.2 provides additional details and context about the model and tests of the model against traditional methods. The main reason this model is able to overcome underestimation of inundation by using only MODIS imagery is because of the temporal nature of the CNN-LSTM. The model is trained not just on concurrent MODIS and Sentinel-1 imagery, but on the 10 weeks preceding the observation as well as topographic variables. This allows the LSTM portion of the model to learn the flood dynamics in Bangladesh and accurately interpolate when a cloud covers a pixel. It does this because the model has learned, for example, what a rising flood wave looks like in time. It has also learned in space that if pixel is flooded and it is at a similar or higher elevation to neighboring pixel , then pixel is also likely flooded. Using the 2022 flood in Sylhet, Bangladesh as a test case, Saunders etΒ al. (2023) show that handcrafted MODIS measures, like MNDWI, indeed fail to predict major floods but that our ML approach with MODIS data replicates the inundated area and temporal flood peak captured by Sentinel-1 with accuracy.
@@
The workflow for our flood model is summarized in Panel B of FigureΒ 3. To detect water in Sentinel-1 data we use a dynamic thresholding algorithm tailored to Bangladesh by Thomas etΒ al. (2023). This produces a binary flood indicator for the 10m pixel. We then upscale the 10m pixel to create a nearly continuous fractional index at 500m resolution, which matches the resolution in MODISβs eight-day composite images. Next we feed this Sentinel-1 derived fractional index, along with hydrologically relevant topographical measures (elevation, slope, and height above nearest drainage), and MODIS 500m eight-day composite data into a CNN-LSTM to predict fractional flood area all the way back to 2002. The data (features) from a single time step are fed into a first CNN which produces a single pixel of spatially contextualized data. This is repeated for each time step to produce a time series of length for that pixel. We then feed a time series of length into the LSTM to produce a single temporal value. The output of the LSTM is then combined with the output of the CNN for time and feed into a second CNN to produce a fractional value for the extent of flooding in a given pixel for a given time step. As with the RF model for rice area, the CNN-LSTM is trained using leave-one-out cross validation, where a time step is withheld for validation one at a time. While Saunders etΒ al. (2023) achieves accuracy in their context, our model is accurate of the time. The result is a time series of the fraction of a 500m pixel that was flooded, covering all of Bangladesh from 2002 to 2022.
3.3 Creation of District-Level EO Dataset